# Privacy Guarantees in Posterior Sampling under Contamination

### v2

@@ B.1.2 Proof for Lemma 7
$\rho_{0,\alpha}(\theta^{*},\theta):=\frac{1}{\alpha}\int\left[\left(\frac{f(x;\theta^{*})}{f(x;\theta)}\right)^{\alpha}-1\right]f(x;\theta^{*})\mathtt{d}x.$
By the assumption that $f(x;\theta)$ has all directional derivatives in $\theta$ and the definition of $\tilde{\bm{\nabla}}$ from Definition 7,
$\left|\left(\frac{f(x;\theta^{*})}{f(x;\theta)}\right)^{\alpha}-1\right|\leq\max_{\theta,\tilde{\theta}\in\bar{S}(c_{r})}\alpha\left(\frac{f(x;\theta^{*})}{f(x;\theta)}\right)^{\alpha-1}\frac{f(x;\theta^{*})}{f(x;\theta)^{2}}\|\tilde{\bm{\nabla}}_{\theta}f(x;\tilde{\theta})\|_{2}\|\theta-\theta^{*}\|_{2}.$
Therefore, if $\|\theta-\theta^{*}\|_{2}\leq c_{r}$, then,
$\rho_{0,\alpha}(\theta^{*},\theta)\leq$ $\,\left(\max_{\theta,\tilde{\theta}\in\bar{S}(c_{r})}\int\left[\frac{f(x;\theta^{*})}{f(x;\theta)}\right]^{\alpha+1}\|\tilde{\bm{\nabla}}_{\theta}f(x;\tilde{\theta})\|_{2}\mathtt{d}x\right)\|\theta-\theta^{*}\|_{2}$
@@ B.1.3 Proof for Lemma 8
    For the case when $r$ is small, let $w_{i}=\theta_{i}-\theta_{i}^{*}$, then $w_{i}\sim\mathcal{N}(\theta^{*}_{i},1)$. Thus $\sum_{i}w_{i}^{2}$ is a non-central $\chi^{2}$ distribution with $d$ degrees of freedom and non-centrality parameter $\lambda=\|\theta^{*}\|_{2}^{2}$ with probability density function given by (Patnaik, 1949)
    $f_{\chi^{2}}(x;d,\lambda)=\frac{1}{2}\exp\left(-\frac{x+\lambda}{2}\right)\left(\frac{x}{\lambda}\right)^{(d-2)/4}I_{d/2-1}(\sqrt{\lambda x})$
@@
    $I_{\nu}(x)=(x/2)^{\nu}\sum_{j=0}^{\infty}\frac{(x^{2}/4)^{j}}{j!\Gamma(\nu+j+1)}$
@@ B.1.4 Proof for Proposition 4
If Assumption 5 holds, then the same assumption holds for the set of density functions $k_{p}(\bm{x};\theta)$ for any $p\in(0,1)$. So it suffices to prove the proposition holds for $f$ given Assumption 5, then the same argument can be used to prove this proposition holds for any $k_{p}$, $p\in(0,1)$.
@@
$A(\psi):=\inf_{\|\theta-\theta^{*}\|_{2}>\psi}h(f(\cdot;\theta),f(\cdot;\theta^{*}))$
@@ B.2 Proofs Related to Lemma 9
Recall that $d(x;\theta):=k_{p}(x;\theta)/k_{p}(x;\theta^{*})$ where $\theta^{*}$ is the true paramter. For $\tau>0$ to be chosen later, let
$\tilde{Z}_{\theta}(\bm{x}):=\max\{-\tau,\log d(\bm{x},\theta)\}$
and
$\nu_{n}(f):=n^{-1/2}\sum_{i=1}^{n}\bm{X}_{i}-\mathbb{E}\bm{X}_{i}.$
@@
    $\mathbb{P}^{*}\left(n^{-1/2}\sum_{i=1}^{n}(X_{i}-\mathbb{E}X_{i})>M\right)\leq\exp\left(-\frac{M^{2}}{2(v+bM/n^{1/2})}\right)$ (28)
@@
Let $\bm{X}_{1},\dots,\bm{X}_{n}$ be i.i.d. random variables following density $k_{p}(\cdot;\theta^{*})$. For any $t>0$, $10<k<$ 1, $M>0$, let
$\Psi(M,t^{2},n):=\frac{M^{2}}{4(8c_{0}t^{2}+M/n^{1/2})}$
@@
$\int_{s(t)/4}^{t}H_{[]}^{1/2}(u,\mathcal{K}_{p}(\Omega_{n}))\mathtt{d}u\leq\frac{\sqrt{3}k^{3/2}M}{2^{10}\exp(\tau/2)}=\frac{\sqrt{3}k^{5/2}n^{1/2}t^{2}}{2^{10}\exp(\tau/2)}$ (29)
@@
$\mathbb{P}^{*}\left(\sup_{\begin{subarray}{c}h(\theta,\theta^{*})\leq t\\ \theta\in\Omega_{n}\end{subarray}}\nu(\tilde{Z}_{\theta})\geq M\right)\leq(1+N(t))\exp(-(1-k)\Psi(M,t^{2},n))$
@@
The proof follows from a similar chaining argument as in Shen and Wong (1994). In this proof only, we will also consider the bracketing entropy computed in $L_{2}$-norm denoted by $H_{[]}(u,\mathcal{F},\|\cdot\|_{2})$. To make further distinction, we will denote the Hellinger bracketing entropy by $H_{[]}(u,\mathcal{F},h)$.
@@
see Lemma 3 of Wong and Shen (1995) or noting that
@@
we can construct a sequence of bracketings on the space $\mathcal{G}$ and $\tilde{\mathcal{Z}}$. Let $t\geq\delta_{0}>\delta_{1}>\cdots>\delta_{N}>0$, and $\mathcal{F}_{0},\dots\mathcal{F}_{N}$ denote the set of Hellinger bracketings of $\mathcal{G}$ with bracket size $\delta_{0},\dots,\delta_{N}$ respectively. We may choose $\delta_{j}$ such that
$\delta_{0}:=\max\{\delta\leq t:H_{[]}(\delta,\mathcal{G},h)\leq\tfrac{k}{4}\Psi(M,t^{2},n)\}$
@@
with $N:=\min\{j:\delta_{j}\leq s\}$. Furthermore, let
$\tilde{\delta}_{0}:=\left(\sup\{\tilde{\delta}\leq 2\sqrt{2}\delta_{0}\exp(\tau/2):H_{[]}(\tilde{\delta},\tilde{\mathcal{Z}},\|\cdot\|_{2})=H_{[]}(\delta_{0},\mathcal{G},h)\}\right)^{-}$
@@
$H_{[]}(\tilde{\delta}_{j},\tilde{\mathcal{Z}},\|\cdot\|_{2})\leq H_{[]}(\delta_{j},\mathcal{G},h)$
@@
For each $\theta\in\Omega_{n}$, there exists a bracket $[\tilde{Z}^{L}_{j,\theta},\tilde{Z}^{U}_{j,\theta}]$ for every $j$ such that $\tilde{Z}^{L}_{j,\theta}\leq\tilde{Z}_{\theta}\leq\tilde{Z}^{U}_{j,\theta}$ and $\|\tilde{Z}^{L}_{j,\theta}-\tilde{Z}^{U}_{j,\theta}\|_{2}\leq\tilde{\delta}_{j}$. Define a sequence of shrinking envelop functions for each $\theta$ by
@@
Then $l_{j}(\theta)\leq\tilde{Z}_{\theta}\leq u_{j}(\theta)$, and $\|u_{j}-l_{j}\|_{2}\leq\tilde{\delta}_{j}$. For notional simplicity, we hide the dependency of $u_{j}(\theta),l_{j}(\theta)$ on $\theta$ and write them directly as $u_{j}$ and $l_{j}$ as functions of $\bm{X}_{i}$. Finally, define sequences $a_{1}>\cdots>a_{N}$, $\eta_{0},\dots,\eta_{N}$ by
@@
$\leq$ $\,\frac{2^{7}}{\sqrt{3k}}\exp(\tau/2)\int_{s/4}^{t}H_{[]}^{1/2}(u,\mathcal{G},h)\mathtt{d}u$
$\leq$ $\,\frac{kM}{8}$
@@
$\leq$ $\,\mathbb{P}_{1}+\mathbb{P}_{2}+\mathbb{P}_{3}+\mathbb{P}_{4}$
@@
$\mathbb{P}_{1}:=$ $\,|\mathcal{F}_{0}|\sup_{h(\theta,\theta^{*})\leq t}\mathbb{P}(\nu_{n}(u_{0})>(1-\tfrac{3k}{8})M)$
$\mathbb{P}_{3}:=$ $\,\sum_{j=0}^{N-1}\mathbb{P}^{*}\left(\sup_{h(\theta,\theta^{*})\leq t}\nu_{n}(\tilde{Z}_{\theta}-u_{j})I_{B_{j}}>\eta_{j}\right)$
@@
Before bounding $\mathbb{P}_{i}$, we can quickly check (31). Note that Hellinger distance has an upper bound of 1, so we may assume $t\leq 1$. Then by definition of $s(t)$, $s(t)\leq t/2\sqrt{2}\leq t/2$, so by $H_{[]}(u,\mathcal{G},h)$ decreasing,
@@
for any $32^{7}c_{0}+16k\leq 2^{14}\cdot 49/$ 3, which can be easily satisfied by a wide range of choices, including the value choice we will use in the proof for main results.
@@
    By the definition of $B_{j}$, on the set $\cup_{i\geq j+1}B_{j}$, we have $-a_{j+1}<l_{j}-u_{j}<u_{j+1}-u_{j}<0$, thus
@@
    $\eta_{j}:=4\tilde{\delta}_{j}k^{-1/2}\left(\sum_{i\leq j+1}H_{[]}(\tilde{\delta}_{i},\tilde{\mathcal{Z}},\|\cdot\|_{2})\right)^{1/2}$
@@
    $\geq$ $\,\frac{3\eta_{j}^{2}}{8\tilde{\delta}_{j}^{2}}\geq 2\sum_{i\leq j+1}\frac{H(\tilde{\delta}_{i},\tilde{\mathcal{Z}},\|\cdot\|_{2})}{k}$
@@
  • For $\mathbb{P}_{3}$ and $\mathbb{P}_{4}$, we try to show that they are both 00.
    $\mathbb{P}_{3}:=\sum_{j=0}^{N-1}\mathbb{P}^{*}\left(\sup_{h(\theta,\theta^{*})\leq t}\nu_{n}(\tilde{Z}_{\theta}-u_{j})I_{B_{j}}>\eta_{j}\right)$
@@
Sum through $\mathbb{P}_{1}$ to $\mathbb{P}_{4}$ completes the proof. ∎
@@
With the same definitions as in Lemma 17, if in addition that
@@
With the same proof as in Lemma 17, but (34) can be instead bounded by
@@
$\leq$ $\,\exp(-(1-k)\Psi(M,t^{2},n))^{2}\left(\sum_{j=0}^{\infty}\exp(-(1-k)4j\Psi(M,t^{2},n)\right)$
@@
By the choice of $\tau$ and $k$ we use later, we see that $t^{2}/2^{10}<s(t)$, thus (20) implies (29), treating $\phi_{n}$ as $t$.If (29) holds for some $t>0$, then it holds for any $\tilde{t}\geq t$. Thus if (20) holds, we may apply the result of Lemma 17 on any $\tilde{t}\geq t$, then
$\mathbb{P}^{*}\left(\sup_{\begin{subarray}{c}h(\theta,\theta^{*})\leq\tilde{t}\\ \theta\in\Omega_{n}\end{subarray}}\nu(\tilde{Z}_{\theta})\geq M(\tilde{t})\right)\leq$ $\,(1+N(\tilde{t}))\exp(-(1-k)\Psi(M,\tilde{t}^{2},n))$
$\leq$ $\,(1+N(t))\exp\left(-\frac{(1-k)k^{2}}{2^{5}c_{0}+4k}n\tilde{t}^{2}\right)$
@@
To derive the choice of constants, we follow the suggestion of Wong and Shen (1995) and choose $k=2/3$, $\exp(\tau/2)=5$, then $c_{0}=25(4-\tau/2)/16\approx 3.74$, $c_{1}=1/12$, $c_{2}=4/27(32c_{0}+8/3)\approx 1.212\times 10^{-3}$ and $c=2^{-7.5}*3^{-1.5}/5\approx 2^{-13}$. Also, for uniform bound, $N\leq\left\lceil 8+\log_{2}(t^{-1})\right\rceil$ and $L=\left\lceil 2\log_{2}(t^{-1})-1\right\rceil$, in which case $(1+L)(1+N)\leq 20\max\{-\log_{2}(t),(\log_{2}(t))^{2}\}$.
For the last result, note that $\exp(-c_{2}nt^{2})\leq\frac{1}{2}$ implies (31), thus by Corollary 18, we can replace $(1+N)$ by $2$ in (37) and get

### v3

@@ B.1.2 Proof for Lemma 6
$\rho_{0,\alpha}(\theta^{*},\theta):=\frac{1}{\alpha}\int\left[\left(\frac{f(\bm{x};\theta^{*})}{f(\bm{x};\theta)}\right)^{\alpha}-1\right]f(\bm{x};\theta^{*})\mathtt{d}\bm{x}.$
By the assumption that $f(\bm{x};\theta)$ has all directional derivatives in $\theta$ and the definition of $\tilde{\bm{\nabla}}$ from Definition 7,
$\left|\left(\frac{f(\bm{x};\theta^{*})}{f(\bm{x};\theta)}\right)^{\alpha}-1\right|\leq\max_{\theta,\tilde{\theta}\in\bar{S}(c_{r})}\alpha\left(\frac{f(\bm{x};\theta^{*})}{f(\bm{x};\theta)}\right)^{\alpha-1}\frac{f(\bm{x};\theta^{*})}{f(\bm{x};\theta)^{2}}\|\tilde{\bm{\nabla}}_{\theta}f(\bm{x};\tilde{\theta})\|_{2}\|\theta-\theta^{*}\|_{2}.$
Therefore, if $\|\theta-\theta^{*}\|_{2}\leq c_{r}$, then
$\rho_{0,\alpha}(\theta^{*},\theta)\leq$ $\,\left(\max_{\theta,\tilde{\theta}\in\bar{S}(c_{r})}\int\left[\frac{f(\bm{x};\theta^{*})}{f(\bm{x};\theta)}\right]^{\alpha+1}\|\tilde{\bm{\nabla}}_{\theta}f(\bm{x};\tilde{\theta})\|_{2}\mathtt{d}\bm{x}\right)\|\theta-\theta^{*}\|_{2}$
@@ B.1.3 Proof for Lemma 7
    For the case when $r$ is small, let $w_{i}=\theta_{i}-\theta_{i}^{*}$. Then $w_{i}\sim\mathcal{N}(\theta^{*}_{i},1)$. Thus $\sum_{i}w_{i}^{2}$ is a non-central $\chi^{2}$ distribution with $d$ degrees of freedom and non-centrality parameter $\lambda=\|\theta^{*}\|_{2}^{2}$, with probability density function given by (Patnaik, 1949)
    $f_{\chi^{2}}(x;d,\lambda)=\frac{1}{2}\exp\left(-\frac{x+\lambda}{2}\right)\left(\frac{x}{\lambda}\right)^{(d-2)/4}I_{d/2-1}(\sqrt{\lambda x}),$
@@
    $I_{\nu}(x)=(x/2)^{\nu}\sum_{j=0}^{\infty}\frac{(x^{2}/4)^{j}}{j!\Gamma(\nu+j+1)}.$
@@ B.1.4 Proof for Proposition 3
If Assumption 5 holds, then the same assumption holds for the set of density functions $k_{p}(\bm{x};\theta)$ for any $p\in(0,1)$. It suffices to prove the proposition holds for $f$ given Assumption 5; the same argument can be used to prove this proposition holds for any $k_{p}$, $p\in(0,1)$.
@@
$A(\psi):=\inf_{\|\theta-\theta^{*}\|_{2}>\psi}h(f(\cdot;\theta),f(\cdot;\theta^{*})),$
@@ B.2 Proofs Related to Lemma 8
Recall that $d(\bm{x};\theta):=k_{p}(\bm{x};\theta)/k_{p}(\bm{x};\theta^{*})$ where $\theta^{*}$ is the true parameter. For $\tau>0$ to be chosen later, let
$\tilde{Z}_{\theta}(\bm{x}):=\max\{-\tau,\log d(\bm{x},\theta)\},$
and for functions $u:\bm{X}\mapsto u(\bm{X})$, define the empirical process
$\nu_{n}(u):=n^{-1/2}\sum_{i=1}^{n}u(\bm{X}_{i})-\mathbb{E}u(\bm{X}_{i}).$
@@
    $\mathbb{P}^{*}\left(n^{-1/2}\sum_{i=1}^{n}(X_{i}-\mathbb{E}X_{i})>M\right)\leq\exp\left(-\frac{M^{2}}{2(v+bM/n^{1/2})}\right),$ (25)
@@
Let $\bm{X}_{1},\dots,\bm{X}_{n}$ be i.i.d. random variables following density $k_{p}(\cdot;\theta^{*})$. For any $t>0$, $0<k<1$, $M>0$, let
$\Psi(M,t^{2},n):=\frac{M^{2}}{4(8c_{0}t^{2}+M/n^{1/2})},$
@@
$\int_{s(t)/4}^{t}H_{[]}^{1/2}(u,\mathcal{K}_{p}(\Omega_{n}))\mathtt{d}u\leq\frac{\sqrt{3}k^{3/2}M}{2^{10}\exp(\tau/2)}=\frac{\sqrt{3}k^{5/2}n^{1/2}t^{2}}{2^{10}\exp(\tau/2)},$ (26)
@@
$\mathbb{P}^{*}\left(\sup_{\begin{subarray}{c}h(\theta,\theta^{*})\leq t\\ \theta\in\Omega_{n}\end{subarray}}\nu(\tilde{Z}_{\theta})\geq M\right)\leq(1+N(t))\exp(-(1-k)\Psi(M,t^{2},n)),$
@@
The proof follows from a similar chaining argument as in Shen and Wong (1994). In this proof only, we will also consider the bracketing entropy computed in $L_{2}$-norm, denoted by $H_{[]}(u,\mathcal{F},\|\cdot\|_{2})$. To make further distinction, we will denote the Hellinger bracketing entropy by $H_{[]}(u,\mathcal{F},h)$.
@@
see Lemma 3 of Wong and Shen (1995) or note that
@@
we can construct a sequence of bracketings on the space $\mathcal{G}$ and $\tilde{\mathcal{Z}}$. Let $t\geq\delta_{0}>\delta_{1}>\cdots>\delta_{N}>0$, and let $\mathcal{F}_{0},\dots\mathcal{F}_{N}$ denote the set of Hellinger bracketings of $\mathcal{G}$ with bracket size $\delta_{0},\dots,\delta_{N}$ respectively. We may choose $\delta_{j}$ such that
$\delta_{0}:=\max\{\delta\leq t:H_{[]}(\delta,\mathcal{G},h)\leq\tfrac{k}{4}\Psi(M,t^{2},n)\},$
@@
with $N:=\min\{j:\delta_{j}\leq s(t)\}$. Furthermore, let
$\tilde{\delta}_{0}:=\left(\sup\{\tilde{\delta}\leq 2\sqrt{2}\delta_{0}\exp(\tau/2):H_{[]}(\tilde{\delta},\tilde{\mathcal{Z}},\|\cdot\|_{2})=H_{[]}(\delta_{0},\mathcal{G},h)\}\right)^{-},$
@@
$H_{[]}(\tilde{\delta}_{j},\tilde{\mathcal{Z}},\|\cdot\|_{2})\leq H_{[]}(\delta_{j},\mathcal{G},h).$
@@
For each $\theta\in\Omega_{n}$, there exists a bracket $\tilde{Z}^{L}_{j,\theta},\tilde{Z}^{U}_{j,\theta}]$ for every $j$ such that $\tilde{Z}^{L}_{j,\theta}\leq\tilde{Z}_{\theta}\leq\tilde{Z}^{U}_{j,\theta}$ and $\|\tilde{Z}^{L}_{j,\theta}-\tilde{Z}^{U}_{j,\theta}\|_{2}\leq\tilde{\delta}_{j}$. Define a sequence of shrinking envelope functions for each $\theta$ by
@@
Then $l_{j}(\theta)\leq\tilde{Z}_{\theta}\leq u_{j}(\theta)$, and $\|u_{j}-l_{j}\|_{2}\leq\tilde{\delta}_{j}$. For notational simplicity, we hide the dependency of $u_{j}(\theta),l_{j}(\theta)$ on $\theta$ and write them directly as $u_{j}$ and $l_{j}$ as functions of $\bm{X}_{i}$. Finally, define sequences $a_{1}>\cdots>a_{N}$, $\eta_{0},\dots,\eta_{N}$ by
@@
$\leq$ $\,\frac{2^{7}}{\sqrt{3k}}\exp(\tau/2)\int_{s/4}^{t}H_{[]}^{1/2}(u,\mathcal{G},h)\mathtt{d}u\leq\frac{kM}{8},$
@@
$\leq$ $\,\mathbb{P}_{1}+\mathbb{P}_{2}+\mathbb{P}_{3}+\mathbb{P}_{4},$
@@
$\mathbb{P}_{1}:=$ $\,|\mathcal{F}_{0}|\sup_{h(\theta,\theta^{*})\leq t}\mathbb{P}(\nu_{n}(u_{0})>(1-\tfrac{3k}{8})M),$
$\mathbb{P}_{2}:=$ $\,\sum_{j=0}^{N-1}\prod_{l_{1}=0}^{j}|\mathcal{F}_{l_{1}}|\prod_{l_{2}=0}^{j+1}|\mathcal{F}_{l_{2}}|\sup_{h(\theta,\theta^{*})\leq t}\mathbb{P}(\nu_{n}(u_{j+1}-u_{j})I_{\cup_{i\geq j+1}B_{i}}>\eta_{j}),$
$\mathbb{P}_{3}:=$ $\,\sum_{j=0}^{N-1}\mathbb{P}^{*}\left(\sup_{h(\theta,\theta^{*})\leq t}\nu_{n}(\tilde{Z}_{\theta}-u_{j})I_{B_{j}}>\eta_{j}\right),$
@@
Before bounding $\mathbb{P}_{i}$, we can quickly check (28). Note that Hellinger distance has an upper bound of 1, so we may assume $t\leq 1$. Then by definition of $s(t)$, $s(t)\leq t/2\sqrt{2}\leq t/2$, so since $H_{[]}(u,\mathcal{G},h)$ is decreasing,
@@
for any $2^{7}c_{0}+16k\leq 2^{14}\cdot 49/3$, which can be easily satisfied by a wide range of choices, including the values we will use in the proof for main results.
@@
    By the definition of $B_{j}$, on the set $\cup_{i\geq j+1}B_{j}$, we have $a_{j+1}<l_{j}-u_{j}<u_{j+1}-u_{j}<0$, thus
@@
    $\eta_{j}:=4\tilde{\delta}_{j}k^{-1/2}\left(\sum_{i\leq j+1}H_{[]}(\tilde{\delta}_{i},\tilde{\mathcal{Z}},\|\cdot\|_{2})\right)^{1/2}.$
@@
    $\geq$ $\,\frac{3\eta_{j}^{2}}{8\tilde{\delta}_{j}^{2}}\geq 2\sum_{i\leq j+1}\frac{H(\tilde{\delta}_{i},\tilde{\mathcal{Z}},\|\cdot\|_{2})}{k},$
@@
  • For $\mathbb{P}_{3}$ and $\mathbb{P}_{4}$, we show that they are both 0.
    $\mathbb{P}_{3}:=\sum_{j=0}^{N-1}\mathbb{P}^{*}\left(\sup_{h(\theta,\theta^{*})\leq t}\nu_{n}(\tilde{Z}_{\theta}-u_{j})I_{B_{j}}>\eta_{j}\right).$
@@
Sum $\mathbb{P}_{1}$ through $\mathbb{P}_{4}$ completes the proof. ∎
@@
With the same definitions as in Lemma 16, if in addition
@@
Following the same proof as in Lemma 16, but (31) can be instead bounded by
@@
$\leq$ $\,\exp(-(1-k)\Psi(M,t^{2},n))^{2}\left(\sum_{j=0}^{\infty}\exp(-(1-k)4j\Psi(M,t^{2},n))\right)$
@@
By the choice of $\tau$ and $k$ we use later, we see that $t^{2}/2^{10}<s(t)$, thus (19) implies (26), treating $\phi_{n}$ as $t$. If (26) holds for some $t>0$, then it holds for any $\tilde{t}\geq t$. Thus, if (19) holds, we may apply the result of Lemma 16 on any $\tilde{t}\geq t$, then
[…]
@@
To derive the choice of constants, we follow the suggestion of Wong and Shen (1995) and choose $k=2/3$, $\exp(\tau/2)=5$, then $c_{0}=25(4-\tau/2)/16\approx 3.74$, $c_{1}=1/12$, $c_{2}=4/27(32c_{0}+8/3)\approx 1.212\times 10^{-3}$ and $c=2^{-7.5}\cdot 3^{-1.5}/5\approx 2^{-13}$. Also, for the uniform bound, $N\leq\left\lceil 8+\log_{2}(t^{-1})\right\rceil$ and $L=\left\lceil 2\log_{2}(t^{-1})-1\right\rceil$, in which case $(1+L)(1+N)\leq 20\max\{-\log_{2}(t),(\log_{2}(t))^{2}\}$.
For the last result, note that $\exp(-c_{2}nt^{2})\leq\frac{1}{2}$ implies (28), thus by Corollary 17, we can replace $(1+N)$ by 22 in (34) and get